in the lab, robyn has two solutions that contain alcohol and is mixing them with each other. Soultion A is 6% alcohol and Solution Bis 20% alcohol. She uses 400 milliliters of Solution A. How many milliters of Solution B does she use, if the resulting mixture is a 12% alcohol solution?
Answer by mananth(16946)

percent ---------------- quantity
Alcohol I 6 ---------------- 400 ml
Alcohol II 20 ---------------- x ml
Mixture 12 ---------------- 400 + x ml

6*400+20x=12( 400 + x )
2400 + 20 x = 4800 + 12 x
20 x + -12 x = 4800 - -2400
8 x = 2400
/ 8
x = 300
300 ml of 20 % Alcohol II